calculate the acceleration in each of the following cases:

(b)a ball starts at rest and after 30 M is velocity has reached 4.8ms^-1
1
Expert's answer
2021-08-29T16:55:45-0400
"v^2=v_0^2+2as,""a=\dfrac{v^2-v_0^2}{2s}=\dfrac{(4.8\ \dfrac{m}{s})^2}{2\cdot30\ m}=0.38\ \dfrac{m}{s^2}."